Week of 170109

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Alberto (monitoring), Charles (databases), Gavin (computing), Jesus (storage), Julia (WLCG), Maarten (SCOD + ALICE + GGUS), Tomas (ATLAS), Vincent (security)
  • remote: Andrew (NLT1), Christian (NDGF), Christoph (CMS), David B (IN2P3), David M (FNAL), Di (TRIUMF), John (RAL), Kyle (OSG), Max (KIT), Onno (NLT1), Victor (JINR), Xin (BNL), Zoltan (LHCb)

Experiments round table:

  • ATLAS reports (raw view) -
    • Tomas: happy new year from ATLAS!
    • Smooth operations over the break. Monte-Carlo, derivations, analysis and small scale reprocessing (express stream) jobs using up to 350K slots, the first three are still running in full speed.
    • One issue with DBRelease misconfiguration (new DBrelease (31.4.1) was missing setup.py file) in HammerCloud jobs causing sites being blacklisted on December 29th, a new DBRelease was built and distributed on CVMFS and sites were manually turned back online in the meantime.
    • Overlay jobs are stressing the frontier servers, trying to throttle the jobs by putting a cap in the # of running (2000).
    • A test is being done this week to evaluate the possibility of running derivations from tape directly, a sample of 100-150TB of AODs will be replicated from DATATAPE to DATADISK at 3 Tier1 sites and we will measure the staging-in times.
    • Preparing for ATLAS Sites Jamboree next week (January 18-20).

  • CMS reports (raw view) -
    • Best wishes for 2017!
    • Run a lot of MC production over Xmas and New Year
      • Frequently utilized 100k CPU cores
      • Also used Tier-0 resources and HLT node (when available)
      • Many thanks to sites and various experts for taking care
    • No major issues to report

  • ALICE -
    • Happy New Year!
    • High to very high activity during the end-of-year break
      • Our thanks go to the sites for their performance and support!
    • A few team tickets were opened, in particular these:
    • On Wed Jan 4 the AliEn central services suffered a scheduled power cut
      • Normally such a cut does not present a problem
      • This time the UPS was exhausted many hours before the power was restored
      • All ALICE grid activity started draining away
      • Some of the hosts needed manual interventions before they would boot
      • The power cut damaged 2 servers, one hosting a critical service
      • That LDAP configuration service was relocated to another host
      • Jobs ran again at CERN by Thu late afternoon
      • The new LDAP service port had to be opened in the campus firewall
      • All was mostly fine again by Fri late morning

  • LHCb reports (raw view) -
    • Zoltan: happy new year from LHCb!
    • Activity
      • very high activity during the Christmas break: running more than 100k jobs (new record for LHCb!)
      • Data reconstruction (proton-ion) almost finished, MC and user analysis.
    • Site Issues
      • T0:
        • NTR
      • T1:
        • Transfer problem from GRIDKA to CBPF (GGUS:125789)
        • RAL: file access issue: User cannot open files (GGUS:125856).

Sites / Services round table:

  • ASGC: Networking device and DPM storage upgrade finished in last downtime. Everything is normal now.
  • BNL: NTR
  • CNAF:
  • EGI:
  • FNAL:
    • FTS, PhEDEx and Xrootd redirector were all upgraded in Dec
  • GridPP:
  • IN2P3: Site availability for ALICE critical tests are not in unknown status for the beginning of the AliEn Central service problem on 4th Jan. It seems it is the same for all sites. Could it be recalculated, thanks? Nothing else to report.
    • Maarten: we will correct for that effect
  • JINR:
    • Our farm still underloaded, 40-50 x 10 cores slots are free for the last few weeks. Site is for CMS only, so a large part of our resources is completely idle.
    • Lost both main and backup links from Dubna to Moscow for full day 05.01.2017, FC intermediate repeaters frozen at -35С.
    • Local network glitch today from 12:00 MSK for around 1 hour, NOC investigating an issue.
  • KISTI:
  • KIT:
    • Unavailability of several SE after failure of disk controllers on 23.12.2016. Provisional solution in place.
    • Network issues at morning of 9.01.2017. Some requests and tests failed as a result. Issues are resolved.
  • NDGF: NTR
  • NL-T1:
    • On new years eve at 23:45 SARA's main mass storage server decided to light some fireworks of its own. A riser card burned down and took two fiber channel cards with it. It took 70 hours before all broken hardware was fixed and tape was accessible again.
    • Network maintenance: one of SARA's public internet uplink fibers has probably been dirty since the datacenter move in October 2016, reducing efficiency to some non-LHCONE/LHCOPN network paths. On December 28, we failed over to our other internet uplink. The dirty fiber will soon be cleaned and recommissioned.
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: Site downtime tomorrow (10th Jan) while we upgrade the Castor NameServer to 2.1.15
  • TRIUMF: Due to a bug in the firmware, one of storage DCS units malfunctioned before Christmas holiday. After working with Vendor and trying to recover the files on that DCS unit, there were still about 2876 files lost.

  • CERN computing services:
    • As mentioned in the ALICE report:
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft reports for the availability of December were sent around today.
    • Final reports for the availability of November are waiting for a GGUS ticket (GGUS:125449). Hopefully the issue will be solved quickly and we should be able to send the final reports by the end of the day.

  • MW Officer: NTR
  • Networks: NTR
  • Security:
    • The incident reported in Dec mainly affected the astrophysics community
    • Two other incidents are under investigation in EGI

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2017-01-09 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback